OpenStudy (anonymous):

Sarah Worker took up golf. She played 36 holes a week for 10 weeks. Use of the course cost $12.00 per round of 18 holes. She paid an 8% sales tax on her equipment expenditures (clubs $275, bag $60, shoes $75, balls and tees $32.50, outfit $35.) How much did her recreation cost Sarah

OpenStudy (rhystic):

so 36/18 = 2 so Sarah has to pay 12 * 2 = 24.00 a week to play. she plays for 10 weeks 24*10 = 240.00. Now add all the equipment expenditures and multiply by 1.08 (due to the 8 percent tax) and add that to 240.

OpenStudy (anonymous):

Sarah's recreational cost is $755.70 Solution: (20 x 12) + .08 x (275 + 60 + 75 + 32.50 + 35) + (275 + 60 + 75 + 32.50 + 35) = $ 755.70